If a red and a blue fair six sided die are rolled what is the probability the result is 8 or divisible by 3? Please try to be as specific as possible with each step. I've how the get the answer but I really don't get how they do it.
Answer #1

1)

total number of outcomes on 2 dice =6*6=36 (as 6 outcome on each dice)

number of outcomes sum can be 8 =5   ({2,6),(3,5),(4,4),(5,3),(6,2)}

hence probability =5/36

2)

number of outcomes sum can be divisible by 3 =12 {(1,2),(2,1),(1,5),(2,4),(3,3).(4,2),(5,1),(3,6),(4,5),(5,4),(6,3),(6,6)}

hence probability =12/36=1/3